OpenStudy (anonymous):

The work of a student trying to solve the equation 4(2x − 1) = 11 + 3x + 5 is shown below: Step 1: 4(2x − 1) = 11 + 3x + 5 Step 2: 8x − 4 = 16 + 3x Step 3: 8x + 3x = 16 − 4 Step 4: 11x = 24 Step 5: x = 2.18 In which step did the student first make an error and what is the correct step? Step 2: 8x − 1 = 16 + 3x Step 2: 8x − 8 = 6 + 3x Step 3: 8x − 3x = 16 + 4 Step 3: 8x + 3x = 16 − 8

OpenStudy (anonymous):

Take a closer look at Step 3... What do you think? ^.~

OpenStudy (anonymous):

I think that 8x + 3x = 16 − 4 should be 8x − 3x = 16 + 4 ??

OpenStudy (anonymous):

CORRECT! xD

OpenStudy (anonymous):

Cewl Thanks again Nina

OpenStudy (anonymous):

^.~ Tha medal?